Ticket Number: HEA-NOC/20070220-6 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: RCSI
Ticket Opened: 20070220 12:42 UTC Problem Start: 20070211 16:47 UTC
Ticket Update: 20070221 15:48 UTC Problem End: 20070227 09:00 UTC
After the RCSI power outage on the 11th the BGP peering was lost between
ar1-cwt and the RCSI router femoral.
Routing did fall back to OSPF, so connectivity was restored when power was
restored, but this is not the preferred routing protocol and it would
appear that the ospf adjacency only formed due to a configuration error.
20070220 12:39 BN Contacted RCSI informing them of the problem. A time
will be agreed for the BGP configuration to be reapplied on their router.
A copy of the proposed configuration was also sent.
20070221 15:46 BN The BGP peering will be reapplied to the RCSI peering
between 08:00 and 09:00 on Tuesday 27th February. While there should be
no outage during this configuration change a short outage to the RCSI
primary connection is possible. This should last no longer than 1 minute.
This ticket can be monitored at http://www.hea.net/tickets/20070220-6